### v1

# Detecting Machine-Generated Music with Explainability – A Challenge and Early Benchmarks

Abstract
Machine-generated music (MGM) has become a groundbreaking innovation with wide-ranging applications, such as music therapy, personalised editing, and creative inspiration within the music industry. However, the unregulated proliferation of MGM presents considerable challenges to the entertainment, education, and arts sectors by potentially undermining the value of high-quality human compositions. Consequently, MGM detection (MGMD) is crucial for preserving the integrity of these fields. Despite its significance, MGMD domain lacks comprehensive benchmark results necessary to drive meaningful progress. To address this gap, we conduct experiments on existing large-scale datasets using a range of foundational models for audio processing, establishing benchmark results tailored to the MGMD task. Our selection includes traditional machine learning models, deep neural networks, Transformer-based architectures, and State Space Models (SSM). Recognising the inherently multimodal nature of music, which integrates both melody and lyrics, we also explore fundamental multimodal models in our experiments. Beyond providing basic binary classification outcomes, we delve deeper into model behaviour using multiple explainable Aritificial Intelligence (XAI) tools, offering insights into their decision-making processes. Our analysis reveals that ResNet18 performs the best according to in-domain and out-of-domain tests. By providing a comprehensive comparison of benchmark results and their interpretability, we propose several directions to inspire future research to develop more robust and effective detection methods for MGM.
Index Terms:
Machine Generated Music Detection, Benchmarking, Explainable AI, Multimodal Models, Baseline
I Introduction
The rise of machine-generated music (MGM), driven by advancements in large language models (LLMs) and exemplified by platforms like MuseNet [1] and AIVA [2], is reshaping the music industry. While this technology offers transformative potential, it also raises critical concerns regarding originality, copyright, and the preservation of artistic value [3]. MGM enhances music production by stimulating creativity, providing compositional frameworks, and suggesting instrumental arrangements [4], yet its capacity for rapid and large-scale output threatens traditional, artist-led creations that rely on time, effort, and individual expression. The growing dominance of Machine-generated styles risks homogenising the music landscape, reducing originality, and influencing public taste through overexposure to algorithmically driven patterns. To address these challenges, effective machine-generated music detection (MGMD) is essential for safeguarding the music community’s creative diversity and fostering a balanced coexistence between human creativity and AI innovation.
Recent studies [5] have critically examined the current landscape of MGMD research, emphasising the urgent need for further progress in this field [6]. The authors identify substantial gaps in the existing literature, particularly the lack of standardised benchmarks, the absence of adequate comparative tools, and the challenges in establishing coherent research directions, especially in the context of explainable artificial intelligence (XAI). This led to the fact that the current MGMD tools are fragmented, often concentrated on narrow areas of investigation. For example, the Spectro-Temporal Tokens Transformer (SpecTTTra) [7] focuses solely on Mel-spectrum features to capture long-range dependencies. In contrast, Afchar [8] employs conventional convolutional methods, placing notable emphasis on the importance of out-of-domain testing and providing insights into output probability for improved interpretability. However, the interpretability remains insufficient, as the approach does not fully incorporate XAI principles. These innovative recent advancements have not to release their models publicly, and their performance comparisons are limited and inadequately explained, highlighting the need for a comprehensive benchmark in this domain.
Recent studies [5] suggest a potential transition from deepfake audio detection to MGMD, given the overlap in the audio features and their shared focus on binary classification. However, music presents unique challenges, particularly with lyrics. Several multimodal models have been developed for music generation, such as lyric-to-melody models [9, 10], which show the efficacy of multimodal feature extraction. Despite these advancements, a comprehensive benchmark for multimodal models in this domain remains lacking.
To address the gaps identified above, we aim to assess the performance of widely used audio classification models as benchmarks, including traditional machine learning models, deep neural network models, Transformer-based models, and state space models (SSM) on the FakeMusicCaps dataset [11]. Subsequently, we conduct out-of-domain testing on the M6 dataset [12] and establish a benchmark for multimodal models. Finally, we leverage and develop new XAI techniques to provide clearer insights into the models’ performance.
Our contributions are:
  • We are the first to provide benchmark results on the MGMD task to the best of our knowledge.
  • We use multimodal models and out-of-domain testing to enhance the benchmark’s scalability and robustness.
  • We introduce new XAI techniques to highlight task challenges and suggest future research directions.
The structure of our paper is as follows. In Section II, we review related works, including fundamental audio processing models, multimodal models, and XAI models. Next, in Section III, we describe the experimental setup and analyse the corresponding results. In Section IV, we apply XAI techniques to explain the performance of the best model. Finally, in Section V, we discuss the challenges encountered in this task and advocate for further research in this area.
II Related work
II-1 Fundamental detection models
Recent studies have reviewed many deepfake audio detection models [13, 14]. Building on these and overlap application with MGMD, the following evaluation categories are proposed: traditional machine learning models, deep neural network models, Transformer-based models [15], State space models (SSM), and multimodal models. The models we selected for evaluation are based on their common use of feature representations in music, particularly the Mel-spectrum.
Traditional machine learning models
Traditional machine learning models are often employed as classifiers. While these models are frequently combined with additional feature processing techniques within a larger pipeline, they can also independently handle the entire task by utilising extracted features to predict labels directly. The (Quadratic) Support Vector Machine ((Q)-SVM) [16] is one of the most commonly utilised models, renowned for its relatively simple parameters. Empirical study [17]has demonstrated the effectiveness of Q-SVM for audio-based classification tasks.
Deep neural network models
Convolutional Neural Networks (CNNs) [18] have demonstrated capability in image-based feature analysis, leading to the development of models built upon them. ResNet18 [19], by leveraging residual connections, addresses the vanishing gradient problem, with applications in deepfake detection shown in work [20]. Additionally, VGG [21] employs deeper layers with small receptive fields, as applied to deepfake audio analysis [22]. To mitigate the inefficiency of such models, SENet [23] introduces squeeze-and-excitation blocks, which refine feature maps [24]. Finally, MobileNet [25] offers a more efficient alternative by utilising depthwise separable convolutions, with its application in synthetic audio detection [26].
On top of models based on CNNs, music typically exhibits a long-form structure that incorporates melody and harmony. To better capture both temporal dependencies and spectral features simultaneously, a hybrid model combining CNNs with LSTM (Long Short-Term Memory) networks [27] has been proposed [28] to enhance the model’s ability to represent the sequential nature of musical data.
Transformer-based models
Transformer-based models [15] have recently surpassed deep neural networks in feature extraction, owing to attention mechanism. The Vision Transformer (ViT) [29], a variant of the Transformer architecture capable of processing images as input, has been successfully applied in audio deepfake detection [30].
SSM
SSM are valued for their ability to model complex dynamic systems and capture temporal dependencies. Mamba [31] employs linear-time sequence modeling to represent dynamic systems and capture long-range dependencies. XLSTM [32] combines SSMs with extended LSTM and an attention-like mechanism to improve modeling of temporal patterns and contextual features. These design choices make both models well-suited for classification tasks.
Multimodal models
Independent of the above choice, there exist multimodal models that focus on extracting audio (melody) and text (lyrics) features, albeit for other applications. Simonetta et al. [33] provide a comprehensive review of multimodal models in music information retrieval, highlighting the modality representation challenges. LLark [34], a multimodal LLM, has been proposed for instruction-following tasks; however, its text processing is designed for instructions rather than lyrics. Despite architectural variations across applications, these models adhere to the principle of extracting features from each modality and subsequently fusing them. As for our benchmark, the simplest approach is enough: features extracted using unimodal models are concatenated before the final classification step, a method known as the early fusion.
II-2 Explainable AI
All the classifiers mentioned above make decisions based on their internal reasoning processes; however, without XAI, the decision-making process remains a black-box, limiting the transparency and interpretability of the results, which reduces their persuasiveness.
Sun et al. [35] recently provided a comprehensive overview of XAI techniques, focusing on their application in the medical domain. Prior studies have also explored the use of XAI in deep learning [36] and acoustic feature analysis for audio deepfake detection [37]. However, XAI techniques specifically tailored to music feature understanding tasks, such as MGMD, remain underexplored. Nevertheless, insights from XAI applications in audio deepfake detection may prove valuable.
There are common foundational XAI techniques: Occlusion Sensitivity [38] evaluates the importance of input regions by masking parts of the input and measuring the resulting changes in the model’s output, but it is time-consuming. In contrast, Integrated Gradients (IG) [39] offer a more efficient way by accumulating gradients along a path from a baseline input to the actual input. With the help of gradients as well, Grad-CAM [40] leverages gradient information to localise critical regions in input images, highlighting areas relevant to the model’s decisions. The basic model of Grad-CAM, Class Activation Mapping (CAM) [41] directly utilises class-specific weights to identify discriminative regions, providing a class-focused perspective. However, CAM requires specific architectural modifications to the model. Unlike the model-specific XAI, model-agnostic model LIME [42] instead approximates local model behaviour using an interpretable surrogate model.
III Experiments
III-A Datasets
For basic evaluation, we selected FakeMusicCaps dataset [11]. This dataset was designed to facilitate the detection of music authorship. It builds upon the MusicCaps dataset [^1] https://paperswithcode.com/dataset/musiccaps, from which human-annotated prompts were extracted and utilised to feed Text-to-Music models for generating MGM with five distinct music generation models. FakeMusicCaps contains 5,521 human-made samples and 27,605 MGM samples, each with a duration of 10 seconds.
For out-of-domain testing and multimodal evaluation, we selected the recently introduced M6 dataset [12], as FakeMusicCaps includes only background music without lyrics. The M6 dataset integrates existing human-made music and generates MGM using multiple generators in conjunction with commercial tools. This comprehensive dataset suits our needs.
We split the dataset into training, validation, and test sets for each type, with proportions of 0.8×0.8, 0.8×0.2, and 0.2, respectively. To facilitate further analysis, all audio files in WAV format were processed with the librosa [43] library to extract the Mel spectrogram representations. To ensure fair experimental comparisons, the same data split was maintained across all experiments. Additionally, all audio samples were resampled to a uniform rate of 16,000 Hz.
III-B Quantitative results
We perform quantitative results to evaluate all models, with detailed hyperparameter choices in supplementary materials.
In-domain test
We conduct the benchmark with the following classification models reviewed in Section II: Q-SVM, ResNet18, VGG, SeNet, MobileNet, CNN with LSTM, and ViT, with their performance metrics shown in “Table I” and AUC-ROC shown in “Figure 1”.
Figure 1: AUC-ROC graph for different models evaluated on FakeMusicCaps.
TABLE I: Training time, Accuracy, and F1 benchmark comparison of models on the FakeMusicCaps dataset
Model Training Time (s) Accuracy F1
Q-SVM[16] 649 .914 .924
ResNet18[19] 96 .924 .924
VGG [21] 268 .564 .408
SeNet [23] 88 .923 .923
MobileNet [25] 99 .968 .968
CNN+LSTM [27] 83 .916 .917
ViT [29] 518 .854 .852
Mamba [31] 22 .690 .682
XLSTM [32] 594 .810 .810
The presented results indicate generally strong performance across the models evaluated. MobileNet demonstrates superior efficacy, achieving an accuracy and F1 measure of 0.968 while maintaining a relatively brief training time of 99 seconds. This highlights its capability as a computationally efficient, yet good performance model for this task. In contrast, the VGG model demonstrates markedly lower performance, achieving an accuracy of 0.564 and an F1 score of 0.408, despite requiring a relatively higher training time of 268 seconds. While Mamba exhibits marginally improved performance, its results remain suboptimal, which may due to the simplicity of its architectural leading to its shorter training time. This discrepancy suggests that the VGG architecture is not well-suited to the dataset or the specific task, although it performs adequately in other image-related tasks. The suboptimal results may also indicate that VGG’s excessive parameter size prevents convergence to an optimal solution. Meanwhile, ResNet18, SeNet, CNN+LSTM, and XLSTM exhibit comparable and relatively strong performance, with ViT demonstrating slightly inferior results, potentially due to the overhead introduced by its attention mechanism. QSVM, while yielding moderate performance, is hindered by high computational complexity due to its quadratic kernel.
Out-of-domian test
Though the performance seems strong, we test the trained models on M6, both their generalised out-of-domain test subset (f) and the other remaining subsets (a-e), with the results shown in TableII.
TABLE II: Out-of-domain model performance on M6. (f) means subset (f) in M6 while (o) means the other remaining ones.
Model Acc(f) F1 (f) Acc(o) F1 (o)
Q-SVM[16] .671 .758 .780 .626
ResNet18[19] .685 .693 .781 .778
VGG[21] .667 .533 .274 .118
SeNet[23] .513 .526 .687 .776
MobileNet[25] .429 .356 .772 .732
CNN+LSTM [27] .392 .316 .776 .737
ViT [29] .623 .630 .762 .754
Mamba [31] .623 .625 .656 .670
XLSTM [32] .405 .422 .710 .712
The results indicate that all models experience a performance drop, highlighting both the poor generalisation capabilities of the benchmarks and the inherent challenges of the task. Specifically, in the out-of-domain test, MobileNet exhibits inferior performance compared to ResNet18. This discrepancy could be attributed to MobileNet’s lightweight architecture, which may reduce its ability to capture complex features. This observation underscores ResNet18 performs the best. Also, it emphasises the need for more structured guidance to enable models to learn truly distinguishable features.
Additionally, performance across the two subsets reveals that some models perform better on subset (f), while others excel on subsets (a-e). This inconsistency suggests that models may be capturing random or spurious features rather than intrinsic ones, highlighting the need for more robust models that can learn fundamental, task-specific representations. Future research should prioritize the design of models capable of effective generalization and focus on extracting intrinsic features. Furthermore, XAI could play a crucial role in guiding the identification of these intrinsic features.
Multimodal model
We also test multimodal models on music with the lyrics subset (bc) of M6. We select an audio processor Wav2Vec 2.0 [44] and a text processor mBERT, a multilingual version of BERT [45] to process the music. The extracted features are concatenated and fed into a Multi-Layer Perceptron (MLP) [46], with optimisation occurring solely during MLP training. Results are presented in Table III.
TABLE III: Performance comparison of multimodal models with others on M6. The other three performances are provided in [12].
Model accuracy F1
SVM .917 .915
ResNet18 .883 .883
ViT .667 .630
Multimodal .975 .975
The results show improved performance with multimodal models, indicating that relying solely on melody may be insufficient. Incorporating additional modalities, such as lyrics, enhances model performance. Future work could explore XAI techniques for multimodal models to show how different modalities contribute to decision-making, thereby improving both interpretability and robustness.
IV Explainable AI
Figure 2: Visualisation of an audio sample using different XAI techniques on the ResNet18 model.
We also explore the explainability of the model in the task of detecting deepfake music. We select ResNet as the representative model for the audio modality due to its strong performance and compatibility with various XAI techniques. We adopt several widely used explainability methods, including IG, LIME, CAM, Grad-CAM, and Occlusion. These methods highlight the regions of input features that the model considers important analyse the decision-making process. We conduct both qualitative and quantitative analyses. Notably, we introduce a novel fidelity experiment in the quantitative analysis by combining multiple XAI techniques. This aims to evaluate the effectiveness and reliability of these explainability techniques in this task through heatmap observations and performance validation based on feature modification.
IV-A Qualitative Analysis
In the qualitative analysis, we investigate the patterns and consistency of different XAI techniques by examining the explainability heatmaps of a single sample. Specifically, we selected a representative deepfake audio sample from the validation set where the model demonstrated high confidence (softmax probability $>$0.9) in its prediction. Using the trained model weights, we generated predictions for this sample and applied various XAI techniques to produce heatmaps, as shown in Figure 2. The heatmaps highlight high-contribution regions critical to the model’s classification. A red-blue transparent colour map is used to visualise the magnitude and polarity of contributions. Red shows positive contributions, while blue represents negative ones. Thresholding is applied to some XAI techniques, focusing on the top 10% of positive and negative contributions. Other regions are made transparent to avoid over-coverage and enhance clarity.
Through observation of the heatmaps, we identify the following patterns. First, the high-contribution regions generated by different XAI techniques show a degree of consistency. For example, in the 0-4 second mid-frequency range and the 6-8 second low-frequency range, IG, CAM, Grad-CAM, and Occlusion all highlight similar high-contribution regions. This suggests that these regions are likely key to the model’s classification decisions. Additionally, the 6-8 second low-frequency range is highlighted as a blue negative contribution region, consistent with the characteristics of the selected audio sample, a real sample correctly predicted as the positive class. In this case, the negative contribution regions in the heatmap represent features that the model identifies as interfering with the classification of the sample as ‘real music’. In other words, the model detects certain uneven frequency distributions and categorises these features as negative contributions correctly. The filtering of potentially disruptive features strengthens the model’s confidence in recognising the sample as real. This behaviour indicates that the model can enhance its recognition of positive samples by effectively ignoring anomalous regions. However, from a perceptual perspective, there is a noticeable break at the 8-second mark in this sample. This might result from a musical variation or pause. Since most XAI models consider this segment negatively correlated with the positive class, it indicates that the model has not fully grasped the essence of musicality as such breaks occur frequently.
Additionally, different XAI techniques show notable differences in their focus on feature breadth and distribution. Grad-CAM and CAM typically concentrate on localised regions of the audio signal. This is likely due to their reliance on the final feature maps of CNNs, which often capture higher-level local information. In contrast, LIME and IG provide finer-grained explanations by focusing on broader regions, covering the entire frequency spectrum. These differences may result from the model’s internal mechanisms and parameter choices. For instance, the baseline used in IG considerably impacts the calculation of feature contributions. In this study, the baseline is set to a zero tensor, which may amplify changes between the input and the baseline. This approach likely guides the model to prioritise regions with greater variations. LIME, however, generates random feature masks to evaluate their influence on the model’s output. While this method enhances input coverage, it introduces uncertainty, particularly in the interpretation of local details, where the stochastic nature of mask generation may compromise the consistency of explanations.
IV-B Quantitative Analysis
To verify the reliability of the XAI results, we design quantitative experiments based on fidelity. The core idea is to mask the regions deemed most important by the model and then reassess its classification performance. This approach quantifies the actual contribution of these regions.
In the single-technique fidelity experiments, we mask the top 10% of high-contribution regions identified by each XAI method, replacing them with zero-filled values. This tests the impact of removing critical regions. Masking is applied to over a thousand validation samples, and metrics such as accuracy, F1, and recall are re-evaluated. As shown in Table IV, all XAI techniques experience a significant drop in performance. For IG, prediction accuracy falls to 50.8%.
TABLE IV: Fidelity experiments with single XAI technology
Visualisation Accuracy F1 Recall
Raw Spectrogram 90.7 92.0 96.2
IG 50.8 64.1 50.8
Grad-CAM 80.5 85.1 98.7
CAM 83.0 87.0 96.6
Occlusion 59.0 73.2 79.0
LIME 56.6 69.8 88.3
Additionally, based on patterns observed in the qualitative analysis, we design a novel multi-technique fidelity experiment. This involves combining the heatmaps generated by 2 to 5 XAI techniques and masking only the overlapping regions. i. e., when the combination size is 2, we mask areas covered by at least two XAI techniques. When the combination size is 3, we mask areas covered by at least three techniques, and so on. We calculate multiple performance indicators to analyse the effect of combining overlapping regions from different XAI techniques, the results are obtained from five independent experiments. as shown in Table V. The “Avg Mask” column represents the average percentage of masked regions. This value decreases significantly as the combination size increases. For example, the average masked area drops from 29.6% at size 2 to just 0.43% at size 5. The accuracy shows the model’s performance on the validation set. Average accuracy improves from 48.6% to 80.0% as the masked area reduces. The p-value quantifies the statistical significance of accuracy changes between successive combination sizes. The improvement between sizes 3 and 4 is significant ($p=0.004$), as is the change between sizes 4 and 5 ($p=0.01$). In contrast, the change between sizes 2 and 3 shows less statistical significance ($p=0.15$). To illustrate the trade-off further, the mask reduction shows the relative decrease in masked area compared to the previous combination size. This reduction reaches 87.4% at size 5. The accuracy change highlights the relative improvement in accuracy. The most notable improvement occurs between sizes 3 and 4, with an increase of 36.7%.
These results demonstrate a clear trade-off between reducing the masked area and maintaining accuracy stability. As the combination size increases, the masked area decreases substantially, from 29.6% to 0.43%. However, accuracy improves gradually and stabilises. This outcome shows that our method effectively identifies the most critical regions for model decision-making. It achieves a balance between minimising the masked area and maintaining stable performance.
TABLE V: Fidelity experiments with multiple XAI technologies
Combination Size Avg Mask Accuracy p-value Mask Reduction Accuracy Change
2 29.6$\pm$7.9 48.6$\pm$4.4 - - -
3 10.2$\pm$2.2 55.6$\pm$5.2 0.15 65.5 14.4
4 3.42$\pm$0.4 76.0$\pm$3.0 0.004 66.5 36.7
5 0.43$\pm$0.3 80.0$\pm$2.2 0.01 87.4 5.3
V Conclusion
In this work, we benchmarked the performance of seven models spanning traditional machine learning, deep neural networks, and Transformer-based methods on the MGMD task, providing a robust set of baselines for future research. Additionally, we explored multimodal approaches, demonstrating their potential in advancing music feature analysis by incorporating complementary modalities such as lyrics.
To assess model robustness and generalisability, we conducted out-of-domain testing and utilised XAI techniques to analyse whether the models learnt intrinsic features. Furthermore, we proposed a novel ensemble XAI strategy to interpret model decisions effectively. Our findings reveal that existing models often fail to leverage fundamental musicological features, such as stops and structural elements, indicating a gap in their ability to capture domain-specific knowledge. This highlights the need for future research to integrate domain-specific knowledge into model architectures, paving the way for more interpretable and robust AI systems in music analysis.